OpenStudy (anonymous):

Tony bought 48 roses for $24. How much do 24 roses cost? $20 $18 $16 $12

OpenStudy (anonymous):

@Frozen15

OpenStudy (anonymous):

12

OpenStudy (frozen15):

@lito_world , no direct answers! Let me explain 12 + 12 = 24 24 + 24 = 48 Ok , so your wondering , WHY did she do this? Well , if you can tell , adding by 12s make the question and the equation match where, the object of this question is already solved/known You already know that 12 + 12 = 24 and 24 + 24 = 48 So we have 48 total of roses Here's a example of what I'm trying to tell/show you Kelly went to the store, there was a total of 48 roses that are $24 Now , Kelly wants to know how much would it costs for a amount of 24 roses? So look at 48 now What is half of that? What number(s) made 48? 24 , yes So with that 24 is half of 48 and 48 roses cost $24 So now , look at 24 roses What is half of that? What number(s) made 24? 12 , yes So then now , you know 12 is half of 24 , cause 12 + 12 =24 You want to know how much would it costs for 24 Think again and read again if 48 roses costs $24 and 12 + 12 = 24 roses Then what is the costs of 24 roses? hmmm?! oooo! I got it! it would costs $12 for a pack of 24 roses , cause , 48 roses costs $24 , which is half of the amount of roses Then i currently know half of 24 is 12 and that is the cost! I currently and would be glad if this helped you , please medal as i DID explain extra facts for you too understand D.$12 is your answer!